Evan Engineering Group receives royalties on a technical manual written by two of its engineers and sold to a publishing company. Royalties are 10% of net sales, receivable on October 1 for sales in January through June and on April 1 for sales in July through December of the prior year. Sales of the manual began in July 2015, and Evan accrued royalty revenue on $33,730 of sales at December 31, 2015.

Evan received royalties of $2,550 on April 1, 2016. On October 1, 2016 Evan received royalties of $3,688. The 2nd half of 2016 sales were estimated to be $43,090

What is Evan's 2016 royalty revenue?
Answer #1

Sales revenue of the year 2015 = $33,730

Accrued royalty revenue on December 31, 2015 = 33,730 x 10%

= $3,373

Evan received royalties of $2,550 on April 1, 2016.

Hence, royalty receivable for the year ended December 31, 2015 = 3,373 - 2,550

= $823

On October 1, 2016, Evan received royalties of $3,688.

Thus, royalty received for the first half of the year 2016 = 3,688 - 823

= $2,865

The 2nd half of 2016 sales were estimated to be $43,090

Hence, royalty for the second half of the year 2016 = 43,090 x 10%

= $4,309

Evan's 2016 royalty revenue = Royalty revenue for the first half + Royalty revenue for the second half

= 2,865 + 4,309

= $7,174

Kindly comment if you need further assistance. Thanks